KTeV Shift Summary - Day 1/14/00


Subject: KTeV Shift Summary - Day 1/14/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:24:03 CST


 
              Shift Summary for day shift, 1/14/00

On Shift: ........

Hours of beam available at all intensities: 8.0 h
Hours of beam available at nominal intensity: 8.0 h
Hours of SHORT_REG data: 0.0 h
Hours of EM reg data: 6.0 h
Hours of special runs (laser scans, muon runs...): 0.3 h

Accounting of lost time & intensity:

        DAQ problems, Access to take regenerator out

========================================================================
Runs ended this shift:
     
       15512 BEAM_MU_SHORT (QKE146-154, QKB164)

Tapes brought to FCC this shift:

      none

========================================================================
Accesses during shift:

       One, to take regenerator/ma out of beam

========================================================================
Overview of shift:

       Special EM regeneration data taking.
       Margaret, Jason and a lot of other people are trying
       to brain storm plane 2 problem.

========================================================================
Description of Shift:

15512 7:00 Muon run with regenerator/MA still in beam line
                      Stoped because of DDD veto 3 problem.
                      The run did stop properly

        ~7:30 Reg/MA are out of beam

15514 8:00 Start first run of electromagnetic regeneration
                      test special run
         8:10 Stoped because DDD veto 2 is raised on ktev2 and
                      was not clear for few minutes. The run stoped fine.

15515 8:20 One more BEAM_EM run. Again DDD veto 2 problem.
                      (after 2 spills). Reboot ktev2.

15516 9:00 One more BEAM_EM run. Still DDD veto 2 problems but
                      they are cleared after few secnods
        10:15 Stop run to increase prescales for triggers 1 & 2
                      Trigger 1 has now prescale factor 4/5.

15517 10:30 One more BEAM_EM run.
        10:45 Again plane 2 is dead, DDD veto 2 is not cleared for
                      few minutes. Stop run. Reboot ktev2 second time.

15518 11:05 One more BEAM_EM run. Plane 2 seems to be better now.
                      It stops working during spills for few seconds
                      but continues running afterwords.

        14:34 Beam is gone...

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shifts



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:24:05 CST